STUDYING THE STEREOCHEMISTRY OF NAPROXEN USING ROTATIONALLY RESOLVED ELEC-TRONIC SPECTROSCOPY. ^{a}

\underline{JUSTIN W. YOUNG}, LEONARDO ALVAREZ-VALTIERRA, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, PA 15260.

Many biochemical processes are stereospecific. An example is the physiological response to a drug that depends on its enantiomeric form. Naproxen is a drug which shows this stereo-specific physiological response. To better understand the stereo specificity of chiral substances, we observed the S_{1} ←S_{0} transitions of R- and S-naproxen in the gas phase using rotationally resolved electronic spectroscopy. The results will be discussed.

aWork supported by NSF (CHE-0615755)